# DiffPatch: Generating Customizable Adversarial Patches using Diffusion Model

### v1

@@ 5 Conclusion
In this work, we proposed a novel diffusion-based adversarial patch generation method DiffPatch to generate naturalistic and customized patches based on a reference image. DiffPatch utilizes image-to-latent mapping, incomplete diffusion optimization, and target mask control to create stylized adversarial patches. Extensive experiments across multiple object detection models validate the effectiveness of DiffPatch. With DiffPatch, we also created a physical-world adversarial dataset, AdvPatch-1K, comprising 1,131 images captured in diverse scenes and conditions. AdvPatch-1K can help develop robust defense methods against adversarial patch attacks.

### v3

@@ 5 Conclusion
In this work, we proposed a novel diffusion-based adversarial patch generation method DiffPatch to generate naturalistic and customized patches based on a reference image. DiffPatch utilizes image-to-latent mapping, incomplete diffusion optimization, and target mask control to create stylized adversarial patches. Extensive experiments across multiple object detection models validate the effectiveness of DiffPatch. With DiffPatch, we also created a physical-world adversarial dataset, AdvT-shirt-1K, comprising 1,131 images captured in diverse scenes and conditions, which validates the feasibility of DiffPatch in physical-world environments.